Johnson Company calculates its allowance for uncollectible accounts as 5% of its ending balance in gross accounts receivable. The allowance for uncollectible accounts had a credit balance of $20,000 at the beginning of 2021. No previously written-off accounts receivable were reinstated during 2021. At 12/31/2021, gross accounts receivable totaled $333,400, and prior to recording the adjusting entry to recognize bad debts expense for 2021, the allowance for uncollectible accounts had a debit balance of 36,600. 


Required:

 1. What was the balance in gross accounts receivable as of 12/31/2020?

 2. What journal entry should Johnson record to recognize bad debt expense for 2021?

 3. Assume Johnson made no other adjustment of the allowance for uncollectible accounts during 2021. Determine the amount of accounts receivable written off during 2021.

 4. If Johnson instead used the direct write-off method, what would bad debt expense be for 2021?
Answer #1

Solution 1:

balance in gross accounts receivable as of 12/31/2020 = allowance for uncollectible accounts balance at beginning of 2021/ 5% = $20000/ 5% = $400,000

Solution 2:

Date Account title Debit Credit
12/31/2021 Bad debt expense ($333400*5% + $36600) 53270
Allowance for uncollectible accounts 53270

Solution 3:

Accounts receivables written off during 2021 = -$36600 - $20000 = $56,600

Soltion 4:

Bad debt expense under direct write off method = accounts receivable written off = $56,600